# Learning in Structured Stackelberg Games

### v2

@@ Abstract
We study structured Stackelberg games, in which both players (the leader and the follower) observe contextual information about the state of the world at time of play. The leader plays against one of a finite number of followers, but the follower’s type is not known until after the game has ended. Importantly, we assume a fixed relationship between the contextual information and the follower’s type, thereby allowing the leader to leverage this additional structure when deciding her strategy. Under this setting, we find that standard learning theoretic measures of complexity do not characterize the difficulty of the leader’s learning task. Instead, we introduce a new notion of dimension, the Stackelberg-Littlestone dimension, which we show characterizes the instance-optimal regret of the leader in the online setting. Based on this, we also provide a provably optimal learning algorithm. We extend our results to the distributional setting, where we use two new notions of dimension, the $\gamma$-Stackelberg-Natarajan dimension and $\gamma$-Stackelberg-Graph dimension. We prove that these control the sample complexity lower and upper bounds respectively, and we design a simple, improper algorithm that achieves the upper bound.
@@ 1 Introduction
Many strategic interactions between a principal and an agent can be cast as Stackelberg games [38]. In a Stackelberg game, the principal (also called the leader) commits to a possibly randomized strategy, before the agent (also called the follower) observes this strategy and best-responds to it. Stackelberg games have had significant real-world impact over the past decade, particularly in security domains [37]. For example, when protecting endangered species [39], a wildlife agency determines ranger patrol locations before poachers select targets. Coast guards decide on shoreline patrol routes in advance of illegal fishers’ actions [17] and, airport security [33] chooses checkpoint screening strategies before smugglers decide which entry points to use.
In many of these settings, the state of the world at time of play can greatly influence the utilities of the players and subsequently the outcome of the game. For example, patrolling some areas of the airport might be deemed of higher- or lower-importance based on the schedule of flights and the number of passengers on board. Similarly, weather patterns may make it harder to reach animals in certain parts of the park, while migration patterns might make others more vulnerable. When contextual information becomes available over time, leveraging it will generally help the leader deploy more effective strategies. This raises a natural question: given that the state of the world might evolve over time, how should the leader in a Stackelberg game design a strategy that adapts to new information coming in over time?
Previous work [27, 13] has formalized such settings as repeated Stackelberg games with side information, where at each interaction both players observe contextual information that affects their utilities. In this setting, followers in different rounds may have different types that determine their payoff for each outcome of the game. Harris et al. [27] show that as long as either the sequence of followers or the sequence of contexts are drawn i.i.d. from a fixed distribution, no-regret learning is possible. That is, the leader is guaranteed to achieve nearly as much utility as the best fixed policy in hindsight. In the fully adversarial regime, however, they show that regret can grow linearly with the time horizon. Interestingly, this lower bound follows from the adversary’s ability to encode an arbitrarily complex relationship between the sequence of contexts and the follower types.
Naturally, there are many cases in which the relationship between the contextual information and the follower about to appear next is not arbitrarily complex, but instead exhibits some structure. In fact, in many settings the contextual information can be predictive of the follower’s type. For example, in wildlife protection, tire tracks might signal the presence of poachers in a particular area of the park, and market demand trends might be indicative of which animals’ poachers are now more likely to appear. Similarly in airport and border security, current affairs might make it more or less likely that certain smuggler groups are active. Motivated by this observation, we study how the leader can learn effectively when there is additional structure between the contextual information and the follower types, thereby circumventing the previous impossibility result.
@@ 1.1 Our Contributions
We revisit the framework of Harris et al. [27] and take a learning theoretic approach to analyzing the leader’s task in predicting the follower type from contexts. Importantly, we find that existing notions of complexity do not capture the difficulty of learning in this setting—we obtain (tight) Stackelberg regret and sample complexity guarantees that are parameterized by new complexity measures and that directly exploit the structure of Stackelberg games. Concretely, our main contributions in this work are as follows.
Instance-optimal Stackelberg regret bounds.
We initiate the study of online learning in Stackelberg games with side information, under the structural assumption that there exists a fixed, unknown mapping between contexts and follower types. In particular, we study how the leader, given access to a predefined hypothesis class, can leverage it to learn. We study the realizable setting, under which the class includes an optimal predictor of the followers’ types. Then the leader’s task is to learn an optimal policy on the fly while achieving as high cumulative utility as possible.
@@
Naturally, the difficulty of the leader’s learning task scales with the complexity of the hypothesis class. In the learning theory literature, the (multiclass) Littlestone dimension [30, 22] is used to characterize the complexity of a class in settings with a finite set of labels (which in structured Stackelberg games corresponds to the follower types). Interestingly, we prove that the Littlestone dimension can give arbitrarily loose regret bounds (Theorem˜3.5). Our construction leverages the fact that the Littlestone dimension is completely blind to the utility space defined by the Stackelberg game, which ultimately determines the performance of the leader.
@@
Based on these insights, we introduce a new notion of complexity, the Stackelberg-Littlestone dimension (Definitions˜3.6 and 3.7, which directly exploits the interaction between the hypothesis class and the utility space of the game. We base our definition on shattered trees, which are the workhorse of classic online learning. The challenge in our setting relies on having a finite set of follower types, yet a continuous utility function, which requires a much more nuanced approach to constructing the trees. We show that the Stackelberg-Littlestone dimension exactly characterizes the difficulty of learning in our setting (Theorem˜3.8), and in particular gives instance-optimal regret bounds. That is, for any hypothesis class and any Stackelberg game, the Stackelberg-Littlestone dimension identifies exactly the optimal regret. The proof of our upper bound Theorem˜3.9 is constructive (Algorithm˜1), which yields a complete characterization of learnability in online structured Stackelberg games.
PAC learning guarantees.
@@
We introduce and formalize the study of distributional learning in structured Stackelberg games (Section˜4). The leader has access to a set of (context, follower type) examples, where the contexts are drawn from some fixed, unknown distribution and the follower types are determined by a true hypothesis in a given hypothesis class. Upon the draw of a new context, the leader’s goal is to deploy a strategy such as to maximize her expected payoff.
@@
We show generalization guarantees with Probably Approximately Correct (PAC)-style bounds. We introduce the notion of $\gamma$-Stackelberg-Natarajan dimension (Definitions˜4.9 and 4.10) and $\gamma$-Stackelberg-Graph dimension (Definitions˜4.12 and 4.13), which respectively control the lower and upper bounds on the sample complexity (Theorems˜4.11 and 4.14). Our definitions are inspired by analogous notions in classification and regression [31, 3, 21], yet we give a tailored approach to what constitutes a “mistake”. This leverages the discreteness of the set of follower types to determine how much hypotheses “differ” in a minmax sense. We give a simple, improper algorithm (Algorithm˜2) that with high probability achieves (almost) optimal expected utility with as many samples as our sample complexity upper bound.
@@
We also analyze a simpler, context-free setting, in which the training set is only over the follower types. Using tools from data-driven algorithm design, we show strong generalization guarantees Theorem˜4.5.
Optimal strategy misprediction bounds.
In the Appendix, we study structured Stackelberg games under a relaxed benchmark that seeks to minimize the total rounds in which the leader does not play the optimal strategy against the realized follower type. We give results for both the online (Appendix˜B) and distributional setting (Section˜B.2), and in both cases, we show stronger guarantees than those achieved by applying existing approaches from learning theory. These results are both conceptually and quantitatively superseded by the ones we present in the main body and we include them for completeness and further insight.
We summarize our results in  Table˜1.
Setting: Online Distributional
Dimension Definitions˜3.6 and 3.7 Definitions˜4.10 and 4.13
Lower Bound Theorem˜3.8 Theorem˜4.11
Upper Bound Theorem˜3.9 Theorem˜4.14
Algorithm Algorithm˜1 Algorithm˜2
Table 1: Summary of our main results.
@@ 1.2 Related Work
Learning in Stackelberg games. The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by Letchford et al. [29] and Blum et al. [16]. These results have been extended by Peng et al. [32], who give faster rates and almost-matching lower bounds. Recently, Bacchiocchi et al. [4] relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy.
@@
Our main focus is on online learning in Stackelberg games. Online Stackelberg games with multiple follower types were introduced in Balcan et al. [7]. The authors show that when a series of unknown attacker types is chosen adversarially, playing the well-known Hedge algorithm  [25] over a carefully selected set of mixed strategies yields a no-regret algorithm.
@@
Our departing point is the work of Harris et al. [27], in which the authors study the problem of online learning in Stackelberg games with side information. The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible. Follow-up work [13] obtains better regret guarantees under bandit feedback. However, when both sequences are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds. In contrast to this line of work, we circumvent this impossibility result by making an additional structural assumption about the relationship between contexts and follower types—namely that it belongs to a class of functions that the leader has access to. Under this assumption, we show that online learning is indeed possible, and our guarantees depend on an appropriate measure of complexity of this function class.
@@
Multiclass learnability. Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types. Then the leader can compute the optimal strategy against the predicted follower. Multiclass prediction has been studied extensively both in the ditributional ([31, 22, 18]) and in the online setting ([22, 26]). Interestingly, we show that learning in Stackelberg games with contexts is at most as hard as multiclass prediction, and in some cases, can be much easier. For example, we give a problem instance the leader can learn to play optimally under a hypothesis class that has infinite Littlestone dimension.
@@
Online Learning beyond Multiclass Prediction. There has been tremendous interest in understanding online learnability in many settings of interest. Notably, in online realizable regression Daskalakis and Golowich [23] designed a proper algorithm that achieved near-optimal cumulative absolute loss with respect to the sequential fat-shattering dimension of the hypothesis class. Our work is more closely related to Attias et al. [3], who introduce the online (resp. $\gamma$-Graph) dimension of a function class and show that it tightly characterizes online regression (resp. distributional learning) in the realizable setting up to a factor of 2. While our setting also considers utility functions with continuous image sets, the dimensions we introduce explicitly exploit the discrete structure of the Stackelberg game that arise from the finite set of follower types. Another recent line of work that captures incentives in online learning is online strategic classification. In this setting, agents can modify their features according to a directed graph over the feature space, before providing these features to the learner. The authors use a new combinatorial measure, the Strategic Littlestone Dimension, to achieve instance optimal mistake bounds. While seemingly similar, this measure of complexity is fundamentally distinct from the one we use to derive our results.
@@
Learning piecewise Lipschitz functions. Given the piecewise linear structure of the leader’s utility function, our work also connects to the line of work on learning functions that are piecewise Lipschitz [8, 36, 10, 11, 9]. We draw from this literature in part of our study that explores a simpler distributional setting without contexts. The well-structured payoff function of the leader against a fixed follower type allows us to adopt tools from Balcan et al. [9] and get better generalization guarantees compared to prior work that studied a similar setting [29].
2 Problem Setting
Notation.
@@ 2 Problem Setting
We use $[N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$. We use $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$. For a given string we use inclusive slice indexing such that $x[i:j]$ are the elements of string $x$ from position $i$ to $j$. Dropping either $i$ or $j$ means that the slice extends from the beginning of the string or to the end, respectively. We use negative indexing to denote the elements of a string starting at the end i.e., $x[-1]$ is the last element of $x$.
@@
We consider a Stackelberg game between a leader and a follower. Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players. The leader plays first and commits to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader. The follower best-responds to the leader’s strategy and the context by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions. [^2] Without loss of generality the follower’s best-response will always be a pure strategy. We define the follower’s best-response as
$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}).$ (1)
@@
Here, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the utility function of follower $f$. In case of ties, we assume that there is a fixed and known ordering over follower actions that determines the follower’s tie-breaking rule.
@@
We allow the follower to be one of $K\in\mathbb{N}_{\geq 1}$ follower types in the set $\{f^{(1)},\dots,f^{(K)}\}$. The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. We assume that the leader knows the set of possible follower types a priori, but not the realized type and aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. By slight abuse of notation, we sometimes use the shorthand $u(\mathbf{z},\mathbf{x},a_{f})=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
@@
For the remainder of the paper and without loss of generality, we will consider $[K]$ to be a set of distinct follower types. We call any two follower types distinct, if the optimal leader’s strategy against them is not the same for every context.
3 Online Learning in Structured Stackelberg Games
In the online setting, the leader is faced with a sequence of followers. In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players. A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best-responds according to Equation˜1. We assume that the leader knows the utilities of all follower types, but the type of follower at time $t$ is not revealed to the leader until after the round is over. We make no assumptions about how the sequence of contexts is chosen, i.e., we allow them to be chosen adversarially and adaptively to the leader’s sequence of play.
3.1 Leader’s objective
Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a set of candidate functions (also called a hypothesis class) mapping contexts to follower types, which the leader uses to learn. We consider the realizable setting, where there exists a true mapping $h^{*}\in\pazocal{H}$ such that $h^{*}(\mathbf{z}_{t})=f_{t}$ for all possible $(\mathbf{z}_{t},f_{t})$ pairs. We use $h^{*}$ as a benchmark for the leader’s performance over time, which we quantify via the notion of contextual Stackelberg regret.

### v3

@@ Abstract
We initiate the study of structured Stackelberg games, a novel form of strategic interaction between a leader and a follower where contextual information can be predictive of the follower’s (unknown) type. Motivated by applications such as security games and AI safety, we show how this additional structure can help the leader learn a utility-maximizing policy in both the online and distributional settings. In the online setting, we first prove that standard learning-theoretic measures of complexity do not characterize the difficulty of the leader’s learning task. Remarkably, we find that there exists a learning-theoretic measure of complexity, analogous to the Littlestone dimension in online classification, that tightly characterizes the leader’s instance-optimal regret. We term this the Stackelberg-Littlestone dimension, and leverage it to provide a provably optimal online learning algorithm. In the distributional setting, we provide analogous results by showing that two new dimensions control the sample complexity upper- and lower-bound.
@@ 1 Introduction
Stackelberg games are the canonical framework for analyzing strategic interactions with commitment, where one player (the leader) commits to a strategy before another (the follower) observes this strategy and best-responds. While initially introduced in the context of market competition [stackelberg1934marktform], Stackelberg games have been applied across a wide range of domains including security, where a security agency commits to a patrol strategy before attackers choose vulnerable targets to attack [Sinha2018StackelbergSG], congestion control, where a city sets tolls before drivers choose routes [swamy2007effectiveness], and AI red-teaming, where AI providers deploy a model before users find adversarial use cases [shivadekar2025red, liu2025chasingmovingtargetsonline].
Contextual Stackelberg games [harrisregret, balcan2025nearlyoptimalbanditlearningstackelberg] are a generalization of this model, in which both players’ payoffs depend on additional side information. This model is motivated by many applications where players have payoff-relevant information about the state of the world at the time of play, which influences their strategies (and subsequently the outcome of the game). For example, a wildlife park’s security might place varying levels of importance on patrolling different areas of the park depending on the time of year and animal migration patterns. Contextual Stackelberg games have also been used to do label-efficient fine-tuning of large language models [wang2026optimal].
Interestingly, learning under this more general framework is not always possible. Specifically, harrisregret study the online setting (where the leader faces a sequence of followers with unknown utilities) and provide no-regret learning guarantees for the leader as long as either the followers or contexts are drawn i.i.d. from a fixed distribution. However, when both contextual information and follower types are chosen adversarially, the worst-case regret can grow linearly with the time horizon. This lower bound follows from a reduction to online classification, where an adversary can encode a mapping from contexts to follower types that is arbitrarily difficult to learn online.
In this work, we ask how broadly this impossibility result applies and investigate the landscape of learnability in this setting. Specifically, we examine learnability in contextual Stackelberg games under an arbitrary, unknown mapping from contexts to follower types. This setting naturally captures domains where contextual data contains some signal that is predictive of the follower about to appear next. In wildlife protection for example, data from motion sensors and camera footage is routinely collected and used to predict the intrusion type [s18051474]. Similarly in AI safety, different types of attack may be more or less viable depending on the domain the model is deployed in. We capture such applications via the structured Stackelberg game model, which we introduce. In this model, we study the following fundamental learnability questions:
(1) In which structured Stackelberg games can the leader learn a utility-maximizing policy?
(2) Are there natural dimension(s) that characterize learnability in this setting?
@@ 1.1 Our contributions
Instance-optimal regret bounds. We initiate the study of online learning in contextual Stackelberg games under an arbitrary, unknown mapping from contexts to follower types. In particular, we study how the leader can learn a utility-maximizing policy, given a hypothesis class rich enough to include this underlying mapping.
@@
Naturally, the difficulty of the leader’s learning task scales with the complexity of the hypothesis class. In the learning theory literature, the (multiclass) Littlestone dimension [littlestone1988learning, pmlr-v19-daniely11a] is used to characterize the complexity of online classification in settings with a finite set of labels (which, in structured Stackelberg games, corresponds to the set of follower types). We prove that the Littlestone dimension can give arbitrarily loose regret bounds in our Stackelberg game setting (Theorem˜3.6). Our construction leverages the fact that the Littlestone dimension is completely blind to the utility space defined by the Stackelberg game, which ultimately determines the performance of the leader.
@@
We introduce a new notion of complexity, the Stackelberg-Littlestone dimension (Definition˜3.9; henceforth SL dimension), designed to capture the joint complexity of the hypothesis class and the utility space of the Stackelberg game. Our definition is based on the notion of shattered trees, which are a classic tool in online learning. However in contrast to previous work, our setting involves a discrete set of follower types, yet a continuous utility function, so constructing the trees requires a more nuanced approach. We prove that the SL dimension tightly characterizes the complexity of learning in structured Stackelberg games. Specifically, we prove that for any hypothesis class and any Stackelberg game, the SL dimension both upper- and lower-bounds the optimal regret for the leader (Theorem˜3.10, Theorem˜3.11). The proof of our upper bound is constructive (Algorithm˜1), which yields a complete characterization of learnability in online structured Stackelberg games.
@@
PAC learning guarantees. We also formalize the study of distributional learning in structured Stackelberg games (Section˜4). In this setting, the leader has access to a set of (context, follower type) examples, where the contexts are drawn from some fixed, unknown distribution and the follower types are determined by some hypothesis in the hypothesis class. Upon the draw of a new context, the leader’s goal is to deploy a strategy such as to maximize her expected utility.
@@
We show generalization guarantees with probably approximately correct (PAC)-style bounds. We introduce the notion of the $\gamma$-SN dimension (Definitions˜4.2 and 4.3) and the $\gamma$-SG dimension (Definitions˜4.5 and 4.6), which control the lower- and upper-bounds on the sample complexity (Theorems˜4.4 and 4.7). Our definitions are inspired by analogous notions in classification and regression [natarajan1989learning, Attias23:regression, daniely2014:optimal], but we rely on a tailored approach to what constitutes a “mistake” in our setting. Our approach leverages the discreteness of the set of follower types to determine how much hypotheses differ in a minmax sense. We give a simple, improper algorithm (Algorithm˜2) that with high probability achieves nearly-optimal expected utility with as many samples as in our sample complexity upper bound.
@@
In the Appendix, we analyze a simpler, context-free offline setting, in which the training set is only over the distribution of follower types. Using tools from data-driven algorithm design, we show strong generalization guarantees (Theorem˜C.8). Moreover, all of our results directly extend to the related problems of learning to bid in auctions with side information and Bayesian persuasion with public and private states. We discuss this in Appendix˜A. Finally, computational considerations are discussed in Appendix˜D.
@@ 1.2 Related work
Learning in Stackelberg games. The problem of learning an optimal strategy for the leader with a single unknown follower type has been studied by letchford2009learning and blum2014learning. These results have been extended by peng2019learning, who give faster rates and almost-matching lower bounds. Recently, bacchiocchi2024sample relax some structural assumptions from previous work, including tie-breaking rules for the follower, and revise worst-case sample complexity results of learning the leader’s optimal mixed strategy.
@@
Online Stackelberg games with multiple follower types were introduced in balcan2015commitment. The authors show that when a series of unknown attacker types is chosen adversarially, playing the Hedge algorithm  [freund1997decision] over a carefully selected set of mixed strategies is sufficient to obtain no-regret.
@@
Our departing point is the work of harrisregret, in which the authors study the problem of online learning in Stackelberg games with side information. The authors show that when either the sequence of follower types or the sequence of contexts are drawn independently from some fixed, unknown distribution, no-regret learning is indeed possible. Follow-up work [balcan2025nearlyoptimalbanditlearningstackelberg] obtains better regret guarantees under bandit feedback. However, when both sequences are chosen adversarially, regret can scale linearly in the worst case, as the problem can be reduced from online learning linear thresholds. We circumvent this impossibility result by making an additional structural assumption about the relationship between contexts and follower types—namely that it belongs to a class of functions that the leader has access to. Under this assumption, we show that online learning is indeed possible, and our guarantees depend on an appropriate measure of complexity of this function class.
wang2026optimal cast the problem of budget-aware supervised fine-tuning of large language models as a contextual Stackelberg game, and provide results for the setting where the leader does not observe the follower’s type by default, but can query it at a cost.
@@
Multiclass learnability. Our work can also be viewed through the lens of multiclass learnability, as it suffices for the leader to (1) learn the hypothesis class of the function mapping the contexts to a (finite) set of follower types, then (2) compute the optimal strategy against the predicted follower. Multiclass prediction has been studied extensively both in the distributional ([natarajan1989learning, pmlr-v19-daniely11a, brukhim2022characterization]) and online ([pmlr-v19-daniely11a, hanneke2023multiclass]) settings. Our results show that just relying on multiclass prediction tools that do not leverage the additional structure of the problem can lead to loose regret guarantees. For example, it is possible to construct a problem instance in which the leader can learn to play optimally under a hypothesis class that has infinite Littlestone dimension.
@@
Learning theory beyond multiclass prediction. In online realizable regression daskalakis2022fast designed a proper algorithm that achieved near-optimal cumulative absolute loss with respect to the sequential fat-shattering dimension of the hypothesis class. Our work is more closely related to Attias23:regression, who introduce the online (resp. $\gamma$-Graph) dimension of a function class and show that it tightly characterizes online regression (resp. distributional learning) in the realizable setting up to a factor of 2. While we also consider utility functions with continuous image sets, the dimensions we introduce explicitly exploit the discrete structure of the Stackelberg game that arises from the finite set of follower types. ahmadi2024strategic introduces a new combinatorial measure, the Strategic Littlestone Dimension, to achieve instance optimal mistake bounds in the online strategic classification setting. While both their measures of complexity and ours characterize learning in various principal-agent settings with commitment, the two are generally incomparable due to the differences in our settings.
@@
Learning piecewise Lipschitz functions. Given the piecewise linear structure of the leader’s utility function, our work also connects to the line of work on learning functions that are piecewise Lipschitz [8555142, Sharma2019LearningPL, balcan2020semi, 10.5555/3540261.3541415, balcan2018general]. Specifically, we draw from this literature in the part of our work that explores a simpler distributional setting without contexts (Section˜C.1). In particular, the well-structured payoff function of the leader against a fixed follower type allows us to adopt tools from balcan2018general and get better generalization guarantees compared to prior work that studies a similar setting [letchford2009learning].
2 Setting and notation
@@
We use $N]:=\{1,2,\ldots,N\}$ to denote the set of natural numbers up to and including $N$ and $\Delta(\pazocal{A})$ to denote the probability simplex over a (finite) set $\pazocal{A}$. For strings, we use inclusive slice indexing such that $x[i:j]$ denotes the entries of $x$ from position $i$ to $j$. Dropping either $i$ or $j$ means that the slice extends from the beginning of the string or to the end, respectively. We also use negative indexing, so that, e.g., x[1]x[-1] is the last element of $x$. All proofs are deferred to the Appendix.
@@
We consider a Stackelberg game between a leader and a follower. Before the game starts, some contextual information $\mathbf{z}\in\pazocal{Z}\subseteq\mathbb{R}^{d}$ is revealed to both players. The leader plays first and commits to a mixed strategy $\mathbf{x}\in\Delta(\pazocal{A})$, where $\pazocal{A}$ is the set of finite actions available to the leader [^1] Our results hold for any strategy set that the leader can optimize over.. The follower best-responds to the leader’s strategy and the context by playing some action $a_{f}\in\pazocal{A}_{f}$, where $\pazocal{A}_{f}$ is the finite set of follower actions. [^2] WLOG the follower’s best-response is a pure strategy. The follower’s best-response is defined as
$b_{f}(\mathbf{z},\mathbf{x})\in\arg\max_{a_{f}\in\pazocal{A}_{f}}\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u_{f}(\mathbf{z},a_{l},a_{f}),$
@@
where, $u_{f}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$ is the follower’s utility function. In case of ties, we assume that the follower uses a fixed, known, and deterministic tie-breaking order over $\pazocal{A}_{f}$.
@@
We allow the follower to be one of $K$ follower types in the set $\{f^{(1)},\dots,f^{(K)}\}$. The type of follower characterizes their utility function, $u_{f^{(i)}}:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. We assume that the leader knows the set of possible follower types a priori [^3] This assumption is for ease of exposition. Our results depend only on the leader knowing the best response function for each follower type., but not the realized type and aims to maximize her own utility, $u:\pazocal{Z}\times\pazocal{A}\times\pazocal{A}_{f}\rightarrow[0,1]$. By slight abuse of notation, we sometimes use the shorthand $u(\mathbf{z},\mathbf{x},a_{f}):=\sum_{a_{l}\in\pazocal{A}}\mathbf{x}[a_{l}]\cdot u(\mathbf{z},a_{l},a_{f}).$
@@
For the remainder of the paper, and without loss of generality, we assume that the follower types in $[K]$ are distinct. We say that two follower types are distinct if there exists some context for which there is no single leader strategy is simultaneously optimal against both types.
Let $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$ be a set of candidate functions (also called a hypothesis class) mapping contexts to follower types, which the leader uses to learn. Throughout, we focus on the realizable setting, i.e., we assume that $H$ is sufficiently rich to contain a perfect mapping $h^{*}\in\pazocal{H}$, where $h^{*}(\mathbf{z}_{t})=f_{t}$ for all $t\in[T]$. We use $h^{*}$ as a benchmark for the leader’s performance. For a single interaction with a follower at context $\mathbf{z}$, we write the leader’s loss when playing strategy $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ as
[…]
We let $\pi_{h^{*}}(\mathbf{z})\in\arg\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z},\mathbf{x},b_{h^{*}(\mathbf{z})}(\mathbf{z},\mathbf{x}))$ denote the optimal policy for the leader.
For a hypothesis class $\pazocal{H}$, we denote its projection on a point as $\pazocal{H}(\mathbf{z})=\{i\in[K]:\exists h\in\pazocal{H}\;\text{s.t.}\;h(\mathbf{z})=i\}$ and on a set as $\pazocal{H}(\{\mathbf{z}_{i}\}_{i\in[n]})=\{(h(\mathbf{z}_{1}),\dots,h(\mathbf{z}_{n})):h\in\pazocal{H}\}$. For $i\in[K]$, we write $\pazocal{H}^{(\mathbf{z}\to i)}=\{h\in\pazocal{H}:h(\mathbf{z})=i\}$.
3 Online learning
In the online setting, the leader is faced with a sequence of followers. In each round $t\in[T]$, a context $\mathbf{z}_{t}\in\pazocal{Z}$ is revealed to both players. A follower of type $f_{t}$ arrives, the leader commits to a mixed strategy $\mathbf{x}_{t}\in\Delta(\pazocal{A})$, and the follower best-responds. We assume that the leader knows the utilities of all follower types, but the type of the follower at time $t$ is not revealed to the leader until after the round is over. We make no assumptions about how the sequence of contexts is chosen, i.e., they may be chosen adversarially.
Next we present the notion of contextual Stackelberg regret, which quantifies the leader’s performance over time.